OpenStudy (anonymous):

Find the equation of the line that passes through (-2,5) and is perpendicular to 2x+2y=18

OpenStudy (vickyisthesmartone123):

how to find the equation of a line perpendicular to a line. Prove that the equation of a line perpendicular to a given line ax + by + c = 0 is bx - ay + λ = 0, where λ is a constant. Let m11 be the slope of the given line ax + by + c = 0 and m22 be the slope of a line perpendicular to the given line. Then, m11 = -abab and m11m22 = -1 ⇒ m22 = -1m11m1 = baba Let c22 be the y-intercept of the required line. Then its equation is y = m22x + c22 ⇒ y = baba x + c22 ⇒ bx - ay + ac22 = 0 ⇒ bx - ay + λ = 0, where λ = ac22 = constant. To get it more clear let us assume that ax + by + c = 0 (b ≠ 0) be the equation of the given straight line. Now convert the ax + by + c = 0 in to slope-intercept form we get, by = - ax - c ⇒ y = - abab x - cbcb Therefore, the slope of the straight line ax + by + c = 0 is (- abab). Let m be the slope of a line which is perpendicular to the line ax + by + c = 0. Then, we must have, m × (- abab) = - 1 ⇒ m = baba Therefore, the equation of a line perpendicular to the line ax + by + c = 0 is y = mx + c ⇒ y = baba x + c ⇒ ay = bx + ac ⇒ bx - ay+ k = 0, where k = ac, is an arbitrary constant.

OpenStudy (vickyisthesmartone123):

This^ explains what to do to solve this! Hopefully it helps! :)

OpenStudy (anonymous):

so k = ac is the final answer? or bx - ay+ k = 0?

OpenStudy (vickyisthesmartone123):

No this is an explanation not the answer. It is an example

OpenStudy (anonymous):

Can you help me solve the problem? Because I get lost every time

OpenStudy (vickyisthesmartone123):

Sorry, but I don't know this...well I do but I get lost too I got this off the internet to maybe it can inform you or help you on how to solve this... :)

OpenStudy (anonymous):

ok

OpenStudy (vickyisthesmartone123):

Sorry @AlixxKK

OpenStudy (anonymous):

it still helped

OpenStudy (vickyisthesmartone123):

:) I'm glad it did!